If the required reserve ratio is 10 percent, currency in circulation is $1,200 billion, checkable deposits are $1,600 billion, and excess reserves total $2,500 billion, Then the money supply is [HINT: As a decimal 0.12345 or percent 12.345%]


1
Expert's answer
2021-05-23T19:12:05-0400

Answer:-


M= M money supply    

MB = monetary base    

C = currency in circulation    

D = checkable deposits    

R = actual reserves    

ER = excess reserves    

q = required reserve ratio  

Money multiplier "= \frac{M}{MB}=\frac{C+D}{R+C}=\frac{C+D}{q×D+ER+C} \\
=\frac{1200+1600}{0.1×1600+2500+1200}=0.73" answer